1. A London dealer bank’s current bid/ask quotes for the U.S. dollar are £0.7720/$ and £0.7980/$. If you want to sell $12,000,000 to that dealer, how many british pounds would you receive?

2. A London dealer bank’s current bid/ask quotes for the U.S. dollar are £0.7720/$ and £0.7980/$. If you want to buy £16,000,000 from that dealer, how many dollars would you pay?

3. How would the NY branch of the London dealer bank in the above question quote its bid and ask prices if it wants to be consistent with the prices of its parent bank?

(Show all your calculations and explain your work. All currency amount answers should be presented with two decimal places, all currency rates with the same number of decimals they are given, and all rates of change and premiums/discounts with four decimal places)
Answer #1

1.Bank’s bid rate is Pound0.7720/$. It means bank will buy dollar at Pounds 0.7720

British pounds received in $12,000,000 = 12,000,000*0.7720 = Pounds 9,264,000

2.The bank’s bid(buying) rate for dollar is Pounds 0.7720

Required Pounds = 16,000,000

Dollar Required = 16,000,000/0.7720

= $20,725,388.60

3. NY Branch’s rate will be 1/0.7980 – 1/0.7720

i.e. $1.2531/Pound and $1.2953/Pound